OpenStudy (patrickcharran97):

Write the slope-intercept form of an equation of the line that passes through the given point and is parallel to the graph of the equation. (4, –3), 3x – 2y = –4

OpenStudy (tracyy):

First of all, let's put the 3x-2y=-4 into y=mx+b form. Subtract 3x. (-2y=-3x-4) Then divide -2 from y. It will look like this:  Simplify and you will get y=3/2 x +2. Since this equation must be parallel to the given, this means that the slope is the same. They will slant the same way. Slope will remain 3/2 x. The y-intercept, however, HAS to be different... In order to get that, you will have to plug in the coordinates (4, -3) into the equation and solve for b, the y-intercept. -3=3/2 (4) +b. Multiply 3/2 with 4 and get.. 6. -3=6 + b Subtract 6 from both sides... -3-6=b And... b=-9. You now have all of the information you need. All you have to do is plug it all in to get your final equation: